# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v2

@@ 3 Our Method: TraCS
3.1 Location Space
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{sta}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$. This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces [9, 10, 37, 16, 29].
@@ 3.2 TraCS-D
Figure 1: Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$. Any other location $\tau_{j}\in\mathcal{S}$ can be represented by direction-distance coordinates $(\varphi,r(\varphi))$ or Cartesian coordinates $(a,b)$.
@@
(a) $\varepsilon=4$
(b) $\varepsilon=6$
Figure 2: Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right. The dominant sector of TraCS-D narrows as $\varepsilon$ increases, leading to a smaller inner-sector (sampling) error. In contrast, the strawman approach has a fixed dominant sector, which is independent of $\varepsilon$ and leads to a large inner-sector error.
3.2.1 Direction Perturbation
@@
Specifically, classical mechanisms like Laplace and Gaussian [13] add unbounded noise, resulting in an output domain of $(-\infty,+\infty)$. This output domain makes them inapplicable to the circular domain $0,2\pi)$. Additionally, these mechanisms (also existing piecewise-based mechanisms) are designed for the distance metric d(y,x)=|yx|d(y,x)=|y-x|, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as the $k$-$R$ mechanism [18], require a finite domain size $k$, making them unsuitable for direct application to the continuous circular domain $0,2\pi)$.
To highlight the necessity and superiority of TraCS-D’s piecewise-based design, we start by presenting a strawman approach that extends the $k$-$R$ mechanism to the circular domain $0,2\pi)$.
Strawman approach: $k$-$R$ + uniform sampling. A former work ATP [37] divides the direction space $0,2\pi)$ into $k$ sectors and applies the $k$-$R$ mechanism. This approach ensures LDP for the $k$ sectors but not for the entire $0,2\pi)$ domain. Even so, we can extend this approach to $0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-$R$.
@@
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within a specific sector. That is, even though the perturbed direction $\varphi^{\prime}$ falls into the same sector as the sensitive direction $\varphi$, the distance between $\varphi$ and $\varphi^{\prime}$ may still be large due to the large sector size. This issue is particularly prominent when $k$ is small, as each sector then spans $2\pi/k$ radians. Uniform sampling within such wide sectors introduces substantial inner-sector errors, which persist regardless of how large the privacy parameter $\varepsilon$ is set. On the other hand, increasing $k$ to reduce sector width impairs the utility of the $k$-$R$ mechanism [32].
@@
We leverage the idea of piecewise-based mechanisms’ design. Specifically, (i) we instantiate a piecewise-based mechanism for the circular domain $0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, which is centered around the sensitive direction with high probability and depends solely on the privacy parameter $\varepsilon$, thus eliminating the need for fixed sectors.
@@ 3.2.2 Distance Perturbation
$|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$ (1)
@@
With the known distance space $0,|\mathcal{D}_{r(\varphi)}|]$ and sensitive distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$. [^2] We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 6 and to align with the circular domain. These two domains are equivalent in implementation. We then design a piecewise-based mechanism for the normalized distance $\overline{r}(\varphi)$ to guarantee LDP on $[0,1)$.
@@
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property [13] ensures it preserves the same privacy level.
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3: Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
3.2.3 Workflow of TraCS-D
@@
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-sensitive reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}$ and $\tau_{i+1}$. For each sensitive location $\tau_{i+1}$, it perturbs the direction (red block) and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9). Line 12 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$ in place, which is then used as the reference for the next iteration ($i+1$). Finally, the algorithm outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
@@ 4  $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
7  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
  $\tau_{i+1}\leftarrow\tau_{i+1}^{\prime}$;
@@ 7  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
8
9return $\mathcal{T}^{\prime}$;
@@ 3.2.4 Analysis of TraCS-D
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if we ignore the complexity of the score function and sampling in the Exponential mechanism [27], NGram [9] has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. L-SRR [29] has $\Theta(mn)$ time complexity, where $m$ is the number of locations, due to the grouping of locations. ATP [37] also has $\Theta(mn)$ time complexity, because of its trajectory merging step.
@@ 3  $\tau_{i}=(a_{i},b_{i})$;
5  $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
@@ 8  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
9
10return $\mathcal{T}^{\prime}$;
@@ 3.3.2 Analysis of TraCS-C
3.4 Rounding to Discrete Space
@@
Heuristic reference location. The possible locations in a discrete space are generally public knowledge, which aids in heuristically selecting a better reference location for TraCS-D. The key observation is that a smaller distance space $\mathcal{D}_{r(\varphi)}$ leads to a smaller distance perturbation error. Therefore, it is beneficial to select a reference location that minimizes the size of $\mathcal{D}_{r(\varphi)}$ to other locations.
@@
Specifically, the direction perturbation from the reference location is likely to be within the dominant sector shown in Figure 2. If a dominant sector contains most of the locations (resulting in smaller rounding error) and is near the boundary of the location space (resulting in smaller distance space), then the corresponding reference location is a good choice.
3.5 Discussions and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, including their privacy, sensitivity to the location space, and extensions to other types of location spaces.
3.5.1 Comparison of Privacy Between TraCS-D and TraCS-C
@@
Although TraCS-D and TraCS-C provide the same level of privacy quantified by $\varepsilon$-LDP—more specifically, location-level (or event-level) LDP as described in surveys [26, 3]—they have different privacy interpretations.
@@
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the usage of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the sensitive distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$. Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction.
@@
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the usage of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive distances $d_{a}$ and $d_{b}$.
3.5.2 Impact of the Size of $\mathcal{S}$

### v3

@@ 3. Our Method: TraCS
3.1. Location Space
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{end}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$. This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/DuHZFCZG23; DBLP:journals/pvldb/Zhang [^0] H23; DBLP:journals/csur/JiangLZZXI21; DBLP:conf/ccs/WangH0QH22).
@@ 3.2. TraCS-D
Figure 1. Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$. Any other location $\tau_{j}\in\mathcal{S}$ can be represented by direction-distance coordinates $(\varphi,r(\varphi))$ or Cartesian coordinates $(a,b)$.
Table 2. Notations used in this paper.
Notation Meaning
$\mathcal{M}_{\circ}$ Direction perturbation mechanism
$\mathcal{M}_{-}$ Distance perturbation mechanism
$\varphi^{\prime}$ Perturbed direction
$r^{\prime}(\varphi)$ Perturbed distance along $\varphi$
(a) $\varepsilon=4$
(b) $\varepsilon=6$
Figure 2. Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right. The dominant sector of TraCS-D narrows as $\varepsilon$ increases, leading to a smaller inner-sector (sampling) error. In contrast, the strawman approach has a fixed dominant sector, which is independent of $\varepsilon$ and leads to a large inner-sector error.
@@
3.2.1. Direction Perturbation
@@
Specifically, classical mechanisms like Laplace and Gaussian (DBLP:journals/fttcs/DworkR14) add unbounded noise, resulting in an output domain of $(-\infty,+\infty)$. This output domain makes them inapplicable to the circular domain $0,2\pi)$. Additionally, these mechanisms are designed for the distance metric d(y,x)=|yx|d(y,x)=|y-x|, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as $k$-$R$ (DBLP:conf/nips/KairouzOV14), assume a finite domain of size $k$ and therefore cannot be directly applied to the continuous circular domain $0,2\pi)$. Recently, OGPM (DBLP:journals/popets/ZhengMH25) proposed a piecewise-based mechanism tailored to circular domains, which inspires our design.
To highlight the necessity and superiority of TraCS-D’s piecewise-based design compared to simpler alternatives, we start by presenting a strawman approach that extends the $k$-$R$ mechanism to the circular domain $0,2\pi)$.
Strawman approach: $k$-$R$ + uniform sampling. A former work ATP (DBLP:journals/pvldb/Zhang [^0] H23) divides the direction space $0,2\pi)$ into $k$ sectors and applies the $k$-$R$ mechanism. This approach ensures LDP for the $k$ sectors but not for the entire $0,2\pi)$ domain. Even so, we can extend this approach to $0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-$R$.
@@
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within a specific sector. That is, even though the perturbed direction $\varphi^{\prime}$ falls into the same sector as the sensitive direction $\varphi$, the distance between $\varphi$ and $\varphi^{\prime}$ may still be large due to the large sector size. This issue is particularly prominent when $k$ is small, as each sector then spans $2\pi/k$ radians. Uniform sampling within such wide sectors introduces substantial inner-sector errors, which persist regardless of how large the privacy parameter $\varepsilon$ is set. On the other hand, increasing $k$ to reduce sector width impairs the utility of the $k$-$R$ mechanism (DBLP:conf/uss/WangBLJ17).
@@
We adapt the design of piecewise-based mechanisms for circular domains from OGPM (DBLP:journals/popets/ZhengMH25). Specifically, (i) we instantiate a piecewise-based mechanism over the circular domain $0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, which is centered around the sensitive direction with high probability and depends solely on the privacy parameter $\varepsilon$, thus eliminating the need for pre-defined fixed sectors.
@@ 3.2.2. Distance Perturbation
(1) $|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$
@@
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3. Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
With the known distance space $0,|\mathcal{D}_{r(\varphi)}|]$ and sensitive distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$. [^6] We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 3 and to align with the circular domain. These two domains are equivalent in implementation. We then employ the piecewise-based mechanism in OGPM (DBLP:journals/popets/ZhengMH25) for the normalized distance $\overline{r}(\varphi)$ to guarantee LDP on $[0,1)$.
@@
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property (DBLP:journals/fttcs/DworkR14) ensures it preserves the same privacy level.
3.2.3. Workflow of TraCS-D
@@
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-sensitive reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}^{\prime}$ and $\tau_{i+1}$ (line 3). For each sensitive location $\tau_{i+1}$, it perturbs the direction (red block) and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9). Line 11 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$, which is then used as the reference for the next iteration ($i+1$). Finally, the algorithm outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
@@ 4  $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
@@
7
8return $\mathcal{T}^{\prime}$;
@@ 3.2.4. Analysis of TraCS-D
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if we ignore the complexity of the score function and sampling in the Exponential mechanism (DBLP:conf/innovations/NissimST12), NGram (DBLP:journals/pvldb/CunninghamCFS21) has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. L-SRR (DBLP:conf/ccs/WangH0QH22) has $\Theta(mn)$ time complexity, where $m$ is the number of locations, due to the grouping of locations. ATP (DBLP:journals/pvldb/Zhang [^0] H23) also has $\Theta(mn)$ time complexity, because of its trajectory merging step.
@@ 3  $\tau_{i}=(a_{i},b_{i})$;
  $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
@@ 7  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
8
9return $\mathcal{T}^{\prime}$;
@@ 3.3.2. Analysis of TraCS-C
3.4. Rounding to Discrete Space
@@
The candidate locations in a discrete space are typically public knowledge, which enables a heuristic selection of a better reference location for TraCS-D. The key observation is that a smaller distance space $\mathcal{D}_{r(\varphi)}$ yields a smaller distance-perturbation error. Therefore, it is preferable to choose a reference location that minimizes the size of $\mathcal{D}_{r(\varphi)}$ with respect to the candidate locations.
@@
Specifically, the direction from the reference location to other locations is often concentrated in a dominant sector, as illustrated in Figure 2. If such a sector covers most candidate locations (thus reducing rounding error) and is close to the boundary of the location space (thus shrinking the distance domain), then the corresponding reference location is a good choice.
3.5. Discussions and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, Euclidean versus spherical geometry, technical considerations in TraCS, and other discussions and extensions.
3.5.1. Comparison of Privacy Between TraCS-D and TraCS-C
@@
Although TraCS-D and TraCS-C provide the same level of privacy quantified by $\varepsilon$-LDP—more specifically, location-level (or event-level) LDP as described in surveys (DBLP:journals/popets/MirandaPascualGPFS23; DBLP:journals/popets/BuchholzAWNK24)—they have different privacy interpretations.
@@
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the use of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the sensitive distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$. Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction.
@@
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the use of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive distances $d_{a}$ and $d_{b}$.
3.5.2. Euclidean Geometry vs Spherical Geometry
TraCS is designed for a rectangular location space $\mathcal{S}\subset\mathbb{R}^{2}$ under Euclidean geometry, where distances are measured by the Euclidean metric. In particular, TraCS-D requires computing the distance spaces $\mathcal{D}_{r(\varphi)}$, which are defined with respect to the Euclidean distance. This choice is made for generality: (i) for many continuous trajectories (e.g. from wearable sensors or indoor devices), Euclidean geometry is a natural choice; and (ii) for city-scale GPS trajectories (e.g. in Chicago and Tokyo), the distortion induced by approximating geographic coordinates as Cartesian coordinates is typically negligible.
Under spherical geometry, distances are measured by the great-circle distance. (i) For country-scale GPS trajectories, standard map projections such as UTM (latlon_utm_converter) can convert GPS coordinates into local Cartesian coordinates, after which Euclidean distance can be used. (ii) For general spherical domains (e.g. the Earth’s surface), we can treat longitude and latitude as two independent coordinates and redesign TraCS-C via independent composition of $(\mathcal{M}_{\circ},\mathcal{M}_{-})$. Specifically, for a location with GPS coordinates $\tau_{i}=(a_{i},b_{i})\in([-\pi,\pi),[-\pi/2,\pi/2])$, i.e. longitude and latitude, we perturb longitude and latitude independently using $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$, respectively. This design does not involve Euclidean distance and aligns with the semantics of longitude and latitude as separate coordinates.
3.5.3. Why Perturb Direction First in TraCS-D
In TraCS-D, we perturb the direction first and then the distance. This order is motivated by two considerations. First, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is location-independent, making it well suited for direct perturbation. Second, the distance space $\mathcal{D}_{r(\varphi)}$ is always defined relative to a specific direction $\varphi$. If we were to perturb the distance before the direction, then after perturbing the direction we would need to redefine the corresponding distance domain using the perturbed direction, which complicates the procedure. Perturbing the direction first avoids this issue and yields a cleaner mechanism design.
3.5.4. Impact of the Size of $\mathcal{S}$